OpenStudy (anonymous):

sin^2x(cot^2x+1)=1. How do I prove the identity?

sam (.sam.):

Use the identity \[\cot^2x+1=\csc^2x\]

OpenStudy (anonymous):

Let me try to show you how I would prove this. \[\cot^s(x)= \cos^2(x)/\sin^2(x)\] This will lead to the following expression: \[\sin^2x((\cos^2x/\sin^2x) + 1) = \sin^2x (( \cos^2x + \sin^2x)/\sin^2x)) = \sin^2x((1/\sin^2x)))=1\] remember that \[\sin^2x+\cos^2x=1 \] the trigonometric pythagoras

OpenStudy (anonymous):

thank you very much. that helped a ton!!

OpenStudy (anonymous):

You are welcome, pardon the slack typesetting, I am still getting used to this.

OpenStudy (anonymous):

no worries.. Im new to this as well so it looks great to me..lol